OpenStudy (anonymous):

Find the expression that represents the length of the hypotenuse of a right triangle whose legs measure m^2-n^2 and 2mn

OpenStudy (triciaal):



OpenStudy (triciaal):

[2 + (m + n)(m -n)]^2 - 2* 2*(m+n)(m-n) = h^2 compare (a + b )^2 - 2ab where a = 2 and b = (m+n)(m-n) h,the hypotenuse, is the squareroot of h^2